OpenStudy (anonymous):

squarre root of 2x+16- square root of x =4

OpenStudy (bahrom7893):

did u mean: \[\sqrt{2x+16}-\sqrt{x} = 4\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

first square both sides. so it should next look like this: 2x+16-x=4^2

OpenStudy (bahrom7893):

u can't do that

OpenStudy (bahrom7893):

theyr not under the same square root

OpenStudy (anonymous):

no? is it becuase they're two different terms?

OpenStudy (anonymous):

oh, right! duhh. basic rules!

OpenStudy (bahrom7893):

okay try other method...

OpenStudy (anonymous):

what method?

OpenStudy (bahrom7893):

id multiply both sides by sqrt(2x+16)+sqrt(x)

OpenStudy (bahrom7893):

to get: 2x+16-x = 4[sqrt(2x+16) + sqrt(x)]

OpenStudy (bahrom7893):

x + 16 = 4(sqrt(2x+16)+sqrtx)

OpenStudy (anonymous):

ok thanks

OpenStudy (bahrom7893):

i honestly dunno what do now.. lol havent done this in a while..

OpenStudy (anonymous):

same here, thats why i messed up. sorry! lemme remember and ill help you in a sec!